OpenStudy (anonymous):

Complete the square: x^2-4x-12

OpenStudy (blurbendy):

(x - 6)(x + 2) = 0 x = 6 x = -2 Do you follow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@nk8522 are you there too?

OpenStudy (anonymous):

@nk8522

OpenStudy (anonymous):

@blurbendy that isn't completing the square, it's just factoring. Completing the square means you want to have it in the form of (x+e)^2 +f. The way you do that is you look at your equation ax^2+bx+c. Take the value of b (in this case -4) and divide it by 2 (in this case you get -2). That's the value of e. So your answer will look like (x-2)^2 + f expand and you get x^2 -4x + 4 How can you get that from your original equation? By adding 4 and subtracting 4 x^2-4x-12 =x^2-4x-12+4-4 =x^2-4x+4-16 (simply rearranging and combining the -12 and the -4) =(x-2)^2-16

OpenStudy (anonymous):

dude thanks! this really helps my algebra takes forever to teach us this. @BluFoot you rock. Best Response.

OpenStudy (anonymous):

*algebra teacher

OpenStudy (anonymous):

@nk8522 are you there?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

yo it worked? good

OpenStudy (anonymous):

so if you want to tag someone you just put "@" and the name of the person and it should open a tag menu of people online

OpenStudy (anonymous):

ya its kinda laggy tho but watevs

OpenStudy (anonymous):

haha ikr xD

OpenStudy (anonymous):

ok is andrew checking this out?

OpenStudy (anonymous):

i sent him the link and anjali as well. i hope he responds ... >.<

OpenStudy (anonymous):

i don't know why he seemed so pissed today.

OpenStudy (anonymous):